OpenStudy (anonymous):

y - 50 = 65/8 (x-10) in standard form ( show work )

OpenStudy (mertsj):

Multiply both sides by 8 first.

OpenStudy (mertsj):

8y-400=65(x+10)

OpenStudy (mertsj):

Remove the parenthesis

OpenStudy (mertsj):

8y-400 = 65x +650

OpenStudy (mertsj):

subtract 65x from both sides

OpenStudy (mertsj):

8y-65x -400 = 650

OpenStudy (mertsj):

Add 400 to both sides

OpenStudy (mertsj):

-65x +8y = 1050

OpenStudy (anonymous):

thank you